OpenStudy (anonymous):

Solve S = 4v2 for v. i need it in steps too please.

OpenStudy (anonymous):

S = 4v^2 v^2 = s/4 v = sqrt(s/4), -sqrt{s/4}

OpenStudy (anonymous):

i'm slightly stupid when it comes to math, so i could really use the help.

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

Didn't we answer it?

OpenStudy (anonymous):

S=4v^2 S/4=v^2 \[v= \pm \sqrt{\S/4}\]

OpenStudy (anonymous):

i have to solve it &explain how i got the answer. &idk how to.

OpenStudy (anonymous):

-divide each side by 4 -square root both sides

OpenStudy (anonymous):

that actually helps me understand more. thanks. (: